Examine the function f (x, y)= xsquare −6xy + 8ysquare −7x + 7y + 14 for extreme values.
1
Expert's answer
2019-03-05T10:49:03-0500

First of all, let us obtain the positions of the critical points by solving the following system of equations:


"\frac{\partial f}{\partial x} =2 x -6 y -7 =0 \\
\frac{\partial f}{\partial y} =-6 x + 16 y +7 =0"

As a result, one can show that (-35/2, -7) is the only critical point. In order to define its type, one should perform the second derivative test, i.e. calculate the value


"A = \frac{\partial^2 f}{\partial x^2} = 2"

and determinant of the Hessian matrix


"D = \begin{vmatrix}
 \frac{\partial^2 f}{\partial x^2} & \frac{\partial^2 f}{\partial x \partial y} \\
 \frac{\partial^2 f}{\partial y \partial x} & \frac{\partial^2 f}{\partial y^2}
\end{vmatrix} = \begin{vmatrix}
 2 & -6 \\
 -6 & 16
\end{vmatrix}
= -4"

at this point. As long as we got D < 0 (the sign of A in this case does not matter), one can conclude that the point (-35/2, -7) is a saddle point.


Answer: there is only one extreme point (-35/2, -7) which is a saddle point.